6/18 Move PR3
- 6/19: DRMI locked.
- 6/21: polarization issue found.
- 6/21: Stop cooler.
- 6/20: Stop cooler
- 6/28: Open chamber
- Long term:
4/22-5/31: Yarm commissioning
5/27-6/30: DRMI commissioning
- 7/1-8/9: FPMI/DRFPMI commissioning
- 8/12-10/4: ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- Please reduce the number or speed of 16kHz channels.
Report of work
- COM: Cancellation of P-polarization by flipped carrier from arm cavity was seen.
- PRXarm helped factor 3 for leaking to P-pol, it means 10% would be 3%.
- CAL: MICH lock trial is going on for calibration.
- CAL: Noise hunting for Xend and found 2 kinds of noise.
- High frequency peak 1-4kHz depending time, and group connection.
Vibration from vacuum pump - > on/off test on Friday.
- Y end work today for the same thing as X end.
- Type A: ETMY was aligned yesterday.
- IP control of ETMY reached to the limit.
- Type B: SRM oplev was tuned and fixed.
- DMG: Fiber channel was installed to secand E18.
- IOO: OMC work today.
- DMG: no maintenance today, but some extra works are planned.